OpenStudy (anonymous):

A package contains 6 candy canes, 2 of which are cracked. If 4 are selected, find the probability of getting no cracked candy canes.

OpenStudy (anonymous):

are they selected and replaced back into the bag? or left out?

OpenStudy (anonymous):

left out

OpenStudy (anonymous):

okay. you will need to do this: there are 6 total, and 2 cracked. so while pulling the first out, the probability of selecting a normal cane would be: \[\frac{ 4 }{ 6 } \] since they are left out, you would only have 5 left, therefore the next probability would be \[\frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

next would be: \[\frac{ 2 }{ 4 }\] followed by: \[\frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

do you follow up to this point?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay great. do you know what to do with the 4 fractions now?